OpenStudy (anonymous):

Find the 9th term of the geometric sequence with a2 = 0.3 and a6 = 0.00003.

OpenStudy (anonymous):

ar = 0.3 ar^5 = 0.00003

OpenStudy (anonymous):

r = 1/10 a = 3 t9 =ar^8 = (3) (0.1)^8 = 0.00000003

OpenStudy (cwrw238):

ar^5 / ar = r^4 = 0,00003/0.3 = 0.0001 r = 0.1 a=3 9th term = ar^8 = 3 * 0.1^8 = 0.00000003